OpenStudy (anonymous):

(2/3)n + 8 = (1/2)n + 2

OpenStudy (anonymous):

(2/3)n + 8 = (1/2)n + 2 2n/3 + 8 = n/2 + 2 2n/3 - n/2 = 2- 8 4n/6 - 3n/6 = -6 n/6 = -6 n=-36

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }n + 8 = \frac{ 1 }{ 2 }n + 2\] I don't understand how you got your answer

OpenStudy (anonymous):

2n/3+8=n/2+2 2n/3-n/2=2-8 4n-3n/6=-6 n/6=-6 n=-36

OpenStudy (anonymous):

2/3n+8=1/2n+2 put the n on the same side so i want to talk the 1/2n to the other so i subtract 1/2n from each side (2/3)n + 8 -1/2n = (1/2)n + 2 -1/2n ====> 2/3n + 8 - 1/2n = 2 then put the numbers on the same side by subtracting an 8 2/3n + 8 - 1/2n -8 = 2 -8 =====> 2n/3 - n/2 = 2- 8 you need a common denominator 2/3 multiply denominator and numerator by 2 ====> 4/6 1/2 multiply denominator and numerator by 3 ====> 3/6 4n/6 - 3n/6 = -6 n/6 = -6 multiply both sides by 6 n=-36

OpenStudy (anonymous):

 Oh my bad diffrent way of doing them then I'm used too.. thank you though